# Generalized hydrodynamics of integrable quantum circuits

### v3

@@ 2.1 The integrable quantum circuit
For a given Hamitlonian, the Suzuki-Trotter decomposition is not unique, but the above discretization preserves integrability [29, 30]. Namely, the Floquet operator (1) features an extensive number of local and quasilocal conserved operators known as charges [29, 30, 31]. For small $\tau$, they can be thought of as a deformation of those of the Hamiltonian (5): for each charge $Q_{k}$, with $[Q_{k},H]=0$, we have two new operators $\tilde{Q}^{\pm}_{k}(\tau)$ with $[\tilde{Q}^{\pm}_{k}(\tau),U(\tau)]=0$. As pointed out in Refs. [29, 31], the charges $\tilde{Q}^{\pm}_{k}(\tau)$ break the single-site translation symmetry $\mathcal{T}$, which maps one onto the other, $\mathcal{T}\tilde{Q}^{\pm}_{k}(\tau)\mathcal{T}^{\dagger}=\tilde{Q}^{\mp}_{k}(\tau)$.
@@ 2.2 The Bethe ansatz solution
and distinguish two cases. First, if $\gamma=i\eta$, with $\eta\in\mathbb{R}$, then $x\in\mathbb{R}$. In this case, we say that the model is in the gapped or massive regime. As it will be clear from the subsequent discussion, the name comes from the fact that the spectrum of $U(\tau)$ can be described in terms of quasiparticles with similar features as those of the XXZ Hamiltonian (5) in the gapped regime $|\Delta|>1$. Second, if $\gamma\in\mathbb{R}$ then $x$ is purely imaginary. In this case, we say that the model is in the gapless or masless regime. The phase diagram as a function of $\Delta$ and $\tau$, which parametrize the gates (4), can be found in Ref. [33]. For any $\Delta$, there is an infinite sequence of transitions as we increase $\tau$. The first phase transition takes place at the value
@@ 2.2.1 The TBA in the gapped regime
In the gapped regime, the string length can take any possible integer value, $n\geq 1$, while the string centers satisfy $\lambda\in[-\pi/2,\pi/2]$. By standard techniques [66], one can then take the thermodydnamic limit of the Bethe equations (9), resulting in an infinite set of coupled integral equations for the distristribution functions $\rho_{n}(\lambda)$, $\rho_{n}^{h}(\lambda)$. Namely, introducing
@@
Note that for $x=0$, the above equations recover the standard TBA description of the homogeneous XXZ Heisenberg chain [66].
@@
The second obervable that we consider is the local current associated with the local spin-$z$ operator. Its form can be inferred from the continuity equation
@@ 3.1 The initial states
Our analysis relies on the possibility to characterise the GGE associated to the chosen initial states. As shown in Ref. [33], this problem can be solved for a large family of states called integrable, generalizing the constructions of the Hamiltonian case [86, 87, 88]. The family of integrable states includes all two-site shift-invariant product states, allowing us to study a wide class of bipartition protocols. For concreteness, we will focus on the following states:
@@ 3.2 GHD equations
$\partial_{t}\rho_{n}(t,x,\lambda)+\partial_{x}(v^{\mathrm{eff}}_{n}(t,x,\lambda)\rho_{n}(t,x,\lambda)),$ (54)
@@ 4.1 The non-interacting points
We begin by discussing the non-interacting points of the model, where we can test our predictions against independent, numerically exact results. These points correspond to the either one of the following choices for the circuit parameters [33]:
@@
The magnetization is vanishing at $\zeta=\pm\infty$, but the profiles is non-trivial inside of the lightcone, signalling the emergence of non-trivial, non-equilibrium steady states. We also note that the magnetization and current profiles display no points of non-analiticities inside of the lightcone. This is expected, since at the non-interacting points we have a single quasi-particle string type [47].
@@
This is a peculiar situation because the Néel and anti-Néel states are related by a translation of a single site. In models in which all charges satisfy single-site translation symmetry, such as the XXZ Heisenberg chain, both states equilibrate to the same GGE [98]. Accordingly, after an initial transient time, there would be no non-trivial dynamics at large space-time scales. This is intuitive when viewing the initial state as created by the addition of a single defect in an homogeneous configuration. Indeed, a local initial perturbation does not typically lead to a change in the macroscopic state of the system, although exceptions exist, see e.g. [99, 100, 101]. On the other hand, the Trotterized XXZ circuit is invariant under translation by two sites, which allows for the possibility that the Néel and anti-Néel states give rise to different GGEs.
@@ Appendix A Numerical solution of the GHD equation
We solve Eq. (A.5) by iteration with initial guess $P^{\prime}_{i}(\lambda)x-E^{\prime}_{i}(\lambda)t$, which converges in the gappless phase, where only finitely many strings are present.

### v4

@@ 2.1 The integrable quantum circuit
For a given Hamiltonian, the Suzuki-Trotter decomposition is not unique, but the above discretization preserves integrability [29, 30]. Namely, the Floquet operator (1) features an extensive number of local and quasilocal conserved operators known as charges [29, 30, 31]. For small $\tau$, they can be thought of as a deformation of those of the Hamiltonian (5): for each charge $Q_{k}$, with $[Q_{k},H]=0$, we have two new operators $\tilde{Q}^{\pm}_{k}(\tau)$ with $[\tilde{Q}^{\pm}_{k}(\tau),U(\tau)]=0$. As pointed out in Refs. [29, 31], the charges $\tilde{Q}^{\pm}_{k}(\tau)$ break the single-site translation symmetry $\mathcal{T}$, which maps one onto the other, $\mathcal{T}\tilde{Q}^{\pm}_{k}(\tau)\mathcal{T}^{\dagger}=\tilde{Q}^{\mp}_{k}(\tau)$.
@@ 2.2 The Bethe ansatz solution
and distinguish two cases. First, if $\gamma=i\eta$, with $\eta\in\mathbb{R}$, then $x\in\mathbb{R}$. In this case, we say that the model is in the gapped or massive regime. As it will be clear from the subsequent discussion, the name comes from the fact that the spectrum of $U(\tau)$ can be described in terms of quasiparticles with similar features as those of the XXZ Hamiltonian (5) in the gapped regime $|\Delta|>1$. Second, if $\gamma\in\mathbb{R}$ then $x$ is purely imaginary. In this case, we say that the model is in the gapless or massless regime. The phase diagram as a function of $\Delta$ and $\tau$, which parametrize the gates (4), can be found in Ref. [33]. For any $\Delta$, there is an infinite sequence of transitions as we increase $\tau$. The first phase transition takes place at the value
@@ 2.2.1 The TBA in the gapped regime
In the gapped regime, the string length can take any possible integer value, $n\geq 1$, while the string centers satisfy $\lambda\in[-\pi/2,\pi/2]$. By standard techniques [66], one can then take the thermodynamic limit of the Bethe equations (9), resulting in an infinite set of coupled integral equations for the distribution functions $\rho_{n}(\lambda)$, $\rho_{n}^{h}(\lambda)$. Namely, introducing
@@
Note that for $x=0$, the above equations recover the standard TBA description of the homogeneous XXZ Heisenberg chain [66]. Eq. (9) is valid for any eigenstate of the system, and the dependence in the eigenstate lies in the set of rapidities $\{p_{j}\}$. In order to derive Eq. (12) one follows the so-called string hypothesis [66], assuming that in the thermodynamic limit the $p$’s arrange into strings of length $n$, whose centers are distributed according to $\rho_{n}(\lambda)$. Eq. (9) then translates into Eq. (12), which relates $\rho_{n}(\lambda)$ to the total density $\rho_{n}^{t}(\lambda)$. Similar to Eq. (9), Eq. (12) holds for any macrostate, and the dependence in a given macrostate is given by an additional relation between $\rho_{n}(\lambda)$ and $\rho_{n}^{t}(\lambda)$ (in other words, fixing $\rho_{n}^{t}/\rho_{n}$ specifies a macrostate uniquely). While $\rho_{n}(\lambda)$ is sufficient to compute expectation values of conserved quantities, $\rho_{n}^{t}(\lambda)$ is needed to compute other quantities, for instance the entropy of the macrostate $\rho_{n}(\lambda)$.
@@
The second observable that we consider is the local current associated with the local spin-$z$ operator. Its form can be inferred from the continuity equation
@@ 3.1 The initial states
Our analysis relies on the possibility to characterize the GGE associated to the chosen initial states. As shown in Ref. [33], this problem can be solved for a large family of states called integrable, generalizing the constructions of the Hamiltonian case [86, 87, 88]. The family of integrable states includes all two-site shift-invariant product states, allowing us to study a wide class of bipartition protocols. For concreteness, we will focus on the following states:
@@ 3.2 GHD equations
$\partial_{t}\rho_{n}(t,x,\lambda)+\partial_{x}(v^{\mathrm{eff}}_{n}(t,x,\lambda)\rho_{n}(t,x,\lambda))$ $=0,$ (54)
@@ 4.1 The non-interacting points
We begin by discussing the non-interacting points of the model, where we can test our predictions against independent, numerically exact results. These points correspond to either one of the following choices of the circuit parameters [33]:
@@
The magnetization is vanishing at $\zeta=\pm\infty$, but the profiles is non-trivial inside of the lightcone, signaling the emergence of non-trivial, non-equilibrium steady states. We also note that the magnetization and current profiles display no points of non-analiticities inside of the lightcone. This is expected, since at the non-interacting points we have a single quasi-particle string type [47].
@@
This is a peculiar situation because the Néel and anti-Néel states are related by a translation of a single site. In models in which all charges satisfy single-site translation symmetry, such as the XXZ Heisenberg chain, both states equilibrate to the same GGE [98]. Accordingly, after an initial transient time, there would be no non-trivial dynamics at large space-time scales. This is intuitive when viewing the initial state as created by the addition of a single defect in an homogeneous configuration. Indeed, a local initial perturbation does not typically lead to a change in the macroscopic state of the system, although exceptions exist, see e.g. [99, 100, 101]. On the other hand, the Trotterized XXZ circuit is invariant under translation by two sites, which allows for the possibility that the Néel and anti-Néel states give rise to different GGEs. It is worth mentioning that one could construct local integrable Hamiltonians breaking the single-site translation symmetry (but preserving two-site shift invariance) for which a similar behavior is observed. However, while this construction would appear rather ad hoc in the continuous-time limit, here it appears naturally as a consequence of increasing the Trotter step. The ability to change the global late-time behavior by a single defect in a circuit setting is also of obvious interest for benchmark experiments in current digital devices. Indeed, approximating the continuous-time evolution of a two-site shift-invariant Hamiltonian would require implementing very deep circuits, even to reach short times, which is unfeasible on current devices. Instead, existing experiments [40, 41, 42] have already shown that integrable circuits can be implemented keeping sufficient coherence over a number of discrete steps that is enough to appreciate the emergence of late-time features.
@@ Appendix A Numerical solution of the GHD equation
We solve Eq. (A.5) by iteration with initial guess $P^{\prime}_{i}(\lambda)x-E^{\prime}_{i}(\lambda)t$, which converges in the gapless phase, where only finitely many strings are present.